Bandar Industries Berhad of Malaysia manufactures sporting equipment. One of the companys products, a football helmet for the North American market, requires a special plastic. During the quarter ending June 30, the company manufactured 3.800 helmets. using 2.204 kilograms of plastic. The plastic cost the company $14,546. According to the standard cost card, each helmet should require 0.51 kilograms of plastic, at a cost of $7.00 per kilogram. Required: 1. What is the standard quantity of kilograms of plastic (SQ) that is allowed to make 3,800 helmets? 2 What is the standard materials cost allowed (SQ SP) to make 3,800 helmets? 3. What is the materials spending variance? 4. What is the materials price variance and the materials quantity variance? (For requlrements 3 and 4, Indicate the effect of each varlance by selecting F for favorable, U for unfavorable, and None for n° effect (Le., zero variance). Input all amounts as posltlve values. Do not round intermediate calculatlons.) 1Standerd quantity of kilograms a 2. Standard cost allowed for actual output allowed 3. Materials spending variance Materials price variance Materials quantity variance

7
Answer #1

Bander Industries answer:

1.Standard quantity of kilograms allowed = 3800 helmets * 0.51 kg = 1938 kg of plastic (SQ)

2. Standard cost allowed for actual output= SQ*SP = 1938 kg * $7 per kg = $13,566

3. Materials spending variance= Standard cost - Actual Cost = $13,566 - $14,546 =$980 Unfavourable

4. Materials price variance= (SP-AP)*AQ = (7-6.60)*2204 kg = $882 Favourable

Actual Price= $14546/2204kg = $6.60 per kg

5. Materials quantity variance = (SQ-AQ)*SP = (1938-2204)*7 = $1862 Unfavourable

Reconciliation :

Material spending variance = Material price variance + Material Quantity variance

Or, $980 Ufavourable = $882 Favourable + $ 1862 Unfavourable

Logistics solutions answer:

1.Standard quantity of labour hours allowed = 175000 items * 0.04 labour hour = 7000 labour hours

2. Standard variable overhead cost allowed = 7000 labour hours * $3.40 per labour hour = $23,800

3. Variable Overhead spending variance= Standard cost - Actual Cost = $23800 - $24790= $990 Unfavourable

4. Variable overhead rate variance= (SR-AR)*Actual Hours= (3.40-3.35)*7400 = $370 favourable

Actual rate= $24790/7400 labour hours= $3.35 per labour hour

5. Variable Overhead Efficiency variance = (SH-AH)SR = (7000-7400)3.40 = $1360 Unfavourable

Reconciliation :

Variable Overhead spending variance = Variable overhead rate variance + Variable Overhead Efficiency variance

Or, $990 Ufavourable = $ 370 Favourable + $ 1360 Unfavourable